OpenStudy (anonymous):

Add, simplify if possible 2/w-7+12/(w-7)^2

OpenStudy (anonymous):

(2/w-7)+[12/(w-7)^2 =(2w-2)/(w-7)^2 or =(2w-2)/(w^2-14w+49)

OpenStudy (anonymous):

if it in this form 2/(w-7) + 12/(w-7)^2 , then [2(w-7) +12] / (w-7)^2 = (2w-2)/(w-7)^2 = 2(w-1)/(w-7)^2